# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

### v2

@@ Abstract
Building Graphical User Interface (GUI) agents is a promising research direction, which simulates human interaction with computers or mobile phones to perform diverse GUI tasks. However, a major challenge in developing generalized GUI agents is the lack of sufficient trajectory data across various operating systems and applications, mainly due to the high cost of manual annotations. In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials. Concretely, we crawl and process online GUI tutorials (such as videos and articles) into GUI agent trajectory data, through which we produce the GUI-Net-1M dataset containing 1M trajectory data across five operating systems and more than 280 applications. We develop the TongUI agent by fine-tuning Qwen2.5-VL-3B/7B models on GUI-Net-1M, which shows remarkable performance improvements on commonly used grounding and navigation benchmarks, outperforming baseline agents about 10% on multiple benchmarks, showing the effectiveness of the GUI-Net-1M dataset and underscoring the significance of our TongUI framework. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
Figure 1: TongUI crawls and processes multimodal web tutorials from diverse sources into GUI trajectories. The collected GUI-Net-1M dataset spans over 280 applications across five operating systems, leading to powerful generalization in a wide range of GUI tasks.
1 Introduction
@@
Graphical User Interface (GUI) agents based on large foundation models are designed to automate tasks on digital devices by emulating human interactions with a variety of applications on different devices [24, 37, 46]. These agents utilize large language models (LLMs) or vision-language models (VLMs) to process visual inputs (screenshots) and textual inputs (accessibility tree and HTML code) and produce corresponding actions, such as clicking buttons, filling forms, and scrolling, to complete GUI tasks [14, 10, 45]. GUI agents significantly enhance human-computer interaction, providing potential applications to various domains, such as software testing, financial services, office assistance, and industrial automation, improving work efficiency and user experience.
@@
Recently, notable efforts have been made for GUI agents by fine-tuning the foundation models [18, 28, 42]. However, collecting sufficient data for fine-tuning is challenging. Most existing approaches rely on either manually annotated interaction trajectories [20] that are high-quality but costly to obtain, or synthetic data generated from large open-source or proprietary LLM/VLMs, which may lack diversity and accuracy [27, 21, 6]. Desirable GUI agents need to perform actions across a wide variety of applications, each with unique interaction sequences, but collecting comprehensive and diverse data across different applications and operating systems remains a major challenge. As a result, the lack of large-scale, diverse, and well-structured GUI data continues to be a key bottleneck in developing robust and generalized GUI agents.
@@
In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials, where ‘Tong’ means generalization in Chinese. Our key observation is that there are readily available multimodal web tutorials on the Internet about how to control computers and smart mobile phones, offering a wealth of information on interacting with various applications and operating systems. These tutorials, in either video or article with screenshots formats, provide detailed step-by-step instructions on interacting with GUI. Compared to the aforementioned manually collected or synthetic data pipeline, online tutorials offer easy accessibility, rich information content, and good quality. Thus, it is natural to build a framework that converts the multimodal web tutorials in diverse formats into task-solving trajectories for GUI agent tuning.
@@
In doing so, we design the TongUI framework composed of four steps: tutorial crawling, tutorial processing, trajectory generation, and data filtering. In the tutorial crawling step, we write some seed tasks and use LLMs to extend them into a wider collection of tasks. The generated tasks serve as keywords for retrieving content from hosts for online GUI tutorials (such as articles from WikiHow and videos from YouTube). The tutorial processing step aims to extract the textual descriptions and screenshots of multimodal tutorials. We first obtain textual descriptions of multimodal tutorials via automatic speech recognition (ASR) or captioning, through which task queries and plans are produced using LLMs on the obtained textual descriptions. Then, we extract salient frames from videos as screenshots of each step, while the images in articles are directly regarded as screenshots. Finally, we leverage one zero-shot GUI agent to automatically recognize trajectories, including reasoning thoughts and actions between two steps. In the data filtering step, we apply a multi-stage pipeline to ensure data quality and GUI relevance, including duplicate tutorial removal, LLM-based content filtering, and trajectory-level filtering using GUI agents. Based on the tutorial collection pipeline, we construct a GUI-Net-1M dataset that contains 1M data across five operating systems with more than 280 applications. Notably, our dataset includes tutorials from different time periods, where evolving layouts still preserve consistent execution logic between similar tasks, helping the model generalize to diverse GUI environments.
@@
In summary, our contributions are three-fold:
(1) We propose the TongUI framework that enables GUI agents to automatically learn from web resources via tutorial crawling, tutorial processing, trajectory generation, data filtering, and agent tuning, leading to better generalization.
(2) We construct the GUI-Net-1M dataset, containing 1M data across five operating systems and over 280 applications. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
(3) We develop the TongUI agent by fine-tuning Qwen2.5-VL models on GUI-Net-1M, achieving improvements on multiple popular benchmarks.
2 Related Work
2.1 GUI Agent
@@
The advancements of LLMs and VLMs accelerate the development of GUI agents. In the early state, only closed-source models are used (e.g., GPT-4, claude 3.5, and ChatGLM). Recently, more and more open-source models are released, such as Show-UI [18], UI-TARS [28], and CogVLM [38]. Grounding and planning are two important capabilities in GUI agents. The grounding capability means whether the GUI could identify correct buttons or regions for operations. Some methods improve the grounding capability by adding more prompts (accessibility trees, HTML, and set-of-mask) [49, 5, 8, 47, 23] and collecting grounding data for fine-tuning [43]. As for the planning capability of GUI agents, existing efforts improve it by prompt engineering [12, 1], supervised fine-tuning [18, 42, 10, 28], or reinforcement learning [13, 26]. In addition, some efforts focus on evaluating the GUI agents from multiple aspects, including action sequence [32, 9], grounding precision [4, 19], trajectory effectiveness [17, 16], and task completion [30, 35, 40]. The evaluating performance in turn guides the research of intelligent agents.
2.2 Agent Data Collection
@@
Compared with data collection for LLMs or VLMs, collecting high-quality data for agents is more challenging, since agents usually require long trajectories that solve complex tasks in different domains (e.g., GUI, multimodal reasoning, embodied AI). Commonly used schemes include human annotation and model synthesis. Human annotation could produce high-quality and diverse trajectory data, but may result in high cost [20]. Some efforts have been made to use zero-shot agents to produce synthetic trajectory data [27, 21, 6], while this may cause out-of-distribution data, leading to inferior results. Similar to other domains, human annotation and model generation are also commonly used to collect trajectories for GUI agents, including the planning, action sequence, and action position [5, 32, 3, 28, 4, 22]. Note that the recent state-of-the-art method UI-TARS [28] has not release its data. In contrast, we will fully release our data.
@@
In addition to the two schemes, considering that there are abundant resources for GUI operations on the Internet, some efforts focus on collecting trajectory data of GUI agents from the Web, to preserve the data quality and reduce costs simultaneously [25, 41]. They collect textual tutorials from the Internet, and convert them into trajectories by using LLMs or exploring in simulators. Unlike them, we directly collect multimodal tutorials from the Internet and propose a multimodal tutorial process pipeline to convert them into trajectories without any simulators. Meanwhile, compared with synthetic data in simulators, our data is closer to tasks in the real world with practical purposes, benefiting to the generalization capability of GUI agents.
Figure 2: Illustration for TongUI model architecture.
3 Method
3.1 Formulation
We formulate GUI tasks as a sequential decision process, and we adopt ReAct [44] as our agent framework. At one time step $i$, given the observation of previous $n$ steps $o_{i-n},\cdots,o_{i-1},o_{i}$ (e.g., a screenshot of GUI), the thoughts $r$ and actions $a$ of previous $n$ steps $(r_{i-n},a_{i-n},\cdots,r_{i-1},a_{i-1})$, and the query $q$, the agent generates a new thought $r_{i}$ and an executable action $a_{i}$ from the action space, such as clicking on a specific UI element, entering text, or scrolling through the interface. Then, executing the action $a_{i}$ leads to a new observation $o_{i+1}$ (e.g., an updated screenshot). This interaction loop continues, with the agent repeatedly observing the environment, selecting actions, and receiving updated observations until either a termination condition is met (e.g., the task is completed or fails) or the maximum number of steps is reached. We parameterize the GUI agent using a vision-language-action model $M_{\theta}$ as shown in Fig. 2,
$r_{i}^{\star},a_{i}^{\star}=\arg\max_{r_{i},a_{i}}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n}\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i}).$ (1)
In this paper, we propose the TongUI framework that aims to tune $M_{\theta}$ by learning from multimodal web tutorials, as shown in Fig. 3. Such tutorials are usually formatted as $\{v,e\}$, where $v$ denotes the visual information (such as images or videos), and $e$ denotes the textual information (such as an introduction to a video). Our goal is to collect diverse multimodal tutorials $\{v,e\}$, and convert them into training data for GUI agents $\{q,o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T}\}$ with the query $q$ and the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps.
3.2 Tutorial Source
To collect multimodal web tutorials, we carefully select multiple data sources via brainstorming, ensuring that we cover a wide range of applications, operating systems, and task types. Concretely, we choose YouTube, Bilibili, TikTok, Wikihow, and Baidu Experience, which host a variety of user-generated content on GUI tasks. YouTube, Bilibili, and TikTok host a variety of video tutorials about GUI tasks. For video tutorials, the visual information $v$ is the video itself, and the textual information $e$ includes the title, the brief introduction to the video, and the caption or audio of the video. Wikihow and Baidu Experience are two widely used platforms that provide step-by-step articles with images in each step across multiple applications and sperating systems. Here, the visual information $v$ is the image sequence among all steps, and the textual information $e$ includes the title and textual content in the article. In this case, we could get diverse multimodal web tutorials.
@@ 3.2 Tutorial Source
3.3 Tutorial Crawling
We use a keyword-based search approach. The search keywords are constructed by combining the name of the applications or website with the task content, that is “app/web + task.” Here, “app/web” refers to the operating system or application being used (e.g., a mobile app or website), while “task” refers to a complete workflow or a set of actions required to accomplish a specific objective (e.g., changing font size in Word, or browsering in Chrome). We identify some task seeds across multiple applications via brainstorming, and expanded them using LLMs for more diverse and relevant tasks.
@@ 3.3 Tutorial Crawling
3.4 Tutorial Process
Textual processing. Given a crawled multimodal tutorial $\{v,e\}$, this process extracts the task $q$ and rough descriptions $\{h_{1},\cdots,h_{T}\}$ in $T$ steps. In video tutorials from Youtube, Tiktok, and Bilibili, the task guidance is usually in the video’s audio information without any textual introductions. To solve this issue, we begin by applying the open-source speech recognition model, Whisper [29], to transcribe the audio streams into the text information $e$. For image-text articles on Wikihow and Baidu Experience, we parse websites based on their article structures for the textual information $e$.
We use LLMs to identify key task-related verbs and nouns, such as “clicking a button”, “filling out a form”, or “selecting a menu”. Then, we leverage both the title and extracted text to classify tutorials into three categories: mobile, desktop, and others. The data with category “others” is discarded. The categories “mobile” and “desktop” allow us to organize the data by device types, ensuring the tasks are appropriately tailored to their respective environments. Finally, we use LLMs to extract the task $q$ and summarize this process into rough descriptions $\{h_{1},\cdots,h_{T}\}$ of $T$ steps from the filtered textual information.
Visual processing. This process extracts the observation $\{o_{1},\cdots,o_{T}\}$ in the $T$ steps. For text-image articles, we could directly obtain the image sequences as the observation $\{o_{1},\cdots,o_{T}\}$ from the crawled data. In practice, there are some noisy data in Baidu Experience and Wikihow, which might be diagrams, comics, or natural images rather than GUI screenshots. To address this issue, we prompt GPT-4o-mini for image classifications, which classify each image as a screenshot or not.
For video tutorials, we need to extract key frames that correspond to critical actions, representing meaningful moments in the task-solving process. We observe that the audio transcript of the videos often contains valuable planning information, such as the description of task steps. Therefore, when audio is available, we first segment the video based on the audio transcript timestamps, treating each segment as a potential task step. We then extract key frames from each of these segments. In the absence of audio, we treat the entire video as a single segment and extract key frames throughout the video. In doing so, we adopt the MOG2 background subtraction algorithm [50, 51, 52] to detect significant changes in GUI interactions.
@@ 3.4 Tutorial Process
After textual and visual processing, tutorials from different sources are structured into a task $q$ with text and image pairs of $T$ steps, denoted as ($\big{(}\{o_{1},h_{1}\},\cdots,\{o_{i},h_{i}\},\cdots,\{o_{T},h_{T}\}\big{)}$, where $o_{i}$ is the image and $h_{i}$ is the rough description in the $i$-th step, respectively.
3.5 Trajectory Generation
@@
This process aims to generate the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps using the task $q$, and image $o_{i}$ and rough description $h_{i}$ in $T$ steps. Concretely, we use a pretrained GUI agent (such as UI-TARS [28]) to generate the thought $r_{i}$ and the action $a_{i}$ using the same action space defined by the GUI agent. For the thought $r_{i}$ and action $a_{i}$, we feed the observation $o_{i}$ with $h_{i}$ as the query to the zero-shot agent. Here, we use $h_{i}$ as the query for agents instead of the task $q$, because $q$ usually contains an abstract goal instead of specific instruction about what to do on the screen. We empirically find that using $h_{i}$ leads to better performance. In this case, each step of text and image pairs can be written as $\{o_{i},r_{i},a_{i}\}$. In practice, the model might fail to generate well-formatted actions. We discard this step and make the next action the beginning of a new trajectory. For example, for a 4-step trajectory which has a failed generation in the second step, we split it into two training trajectories: ($(\{o_{1},r_{1},a_{1}\})$ and ($(\{o_{3},r_{3},a_{3}\},\{o_{4},r_{4},a_{4}\})$. Finally, we combine the trajectory with the task $q$ as $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$. We collect these trajectories into a GUI-Net-1M dataset that contains 1M data. These data is on five operating systems with more than 280 applications. After that, we use GUI-Net-1M for supervised fine-tuning (SFT).
3.6 Data Filtering
@@
We apply a multi-stage data filtering pipeline for data quality, as shown in Fig. 3. (1) Duplicate Tutorial Removal. To eliminate redundancy of crawling, we directly remove exact duplicate tutorials based on their unique identifiers (e.g., video IDs and URLs). This step ensures that each retained tutorial is unique, avoiding repetition and simplifying downstream processing. (2) Content Filtering. After tutorial processing, we employ an LLM to semantically filter out tutorials that are irrelevant to the GUI tasks. Concretely, based on the tutorial titles and audio transcriptions, the LLM evaluates whether the content is centered around GUI interactions. (3) Trajectory Filtering. In the trajectory generation stage, since we use UI-TARS [28], if a screenshot corresponds to a step unrelated to GUI interaction, the agent typically predicts actions like wait or call_user. We use this as a signal to discard such observations. Then, we feed the screenshot and trajectory of a task into a Qwen2.5-VL-7B model, and prompt it to filter out low-quality data.
3.7 Agent Tuning
@@
Given a data point $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$ of $T$-step, we train a VLM $M_{\theta}$ using the cross-entropy loss,
$\min\mathbb{E}_{(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})\sim\mathbb{D}}\big{[}-\sum_{i=1}^{T}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n},\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i})\big{]},$

### v3

@@ Abstract
Building Graphical User Interface (GUI) agents is a promising research direction, which simulates human interaction with computers or mobile phones to perform diverse GUI tasks. However, a major challenge in developing generalized GUI agents is the lack of sufficient trajectory data across various operating systems and applications, mainly due to the high cost of manual annotations. In this paper, we propose the TongUI framework that transforms millions of multimodal web tutorials into GUI trajectories for generalized GUI agents. Concretely, we crawl GUI videos and articles from the Internet and process them into GUI agent trajectory data. Based on this, we construct the GUI-Net-1M dataset, which contains 1 million trajectories across five operating systems and over 280 applications. To the best of our knowledge, this is the largest open-source GUI trajectory dataset. We develop the TongUI agent by fine-tuning Qwen2.5-VL-3B/7B/32B models on GUI-Net-1M, which shows consistent performance improvements on commonly used grounding and navigation benchmarks, outperforming baseline agents by 10% on multiple benchmarks, showing the effectiveness of the GUI-Net-1M dataset and underscoring the significance of our TongUI framework.
Projecthttps://computer-use-agents.github.io/tongui/
Codehttps://github.com/TongUI-agent/TongUI-agent
Datasethttps://huggingface.co/datasets/Bofeee5675/ GUI-Net-1M
Introduction
@@
Graphical User Interface (GUI) agents based on large foundation models are designed to automate tasks on digital devices by emulating human interactions with a variety of operating systems and applications (nguyen2024gui; wang2024gui; zhang2024large). These agents utilize large language models (LLMs) or vision-language models (VLMs) to process visual inputs (screenshots) and textual inputs (accessibility tree and HTML code), and produce corresponding actions, such as clicking buttons, filling forms, and scrolling, to complete GUI tasks (lee2023explore; hong2024cogagent; you2024ferret). GUI agents significantly enhance human-computer interaction, providing potential applications to various domains, such as software testing, financial services, office assistance, and industrial automation, improving work efficiency and user experience.
@@
Recently, notable efforts have been made for GUI agents by fine-tuning the foundation models (lin2024showui; qin2025ui; xu2024aguvis). However, collecting sufficient trajectory data for fine-tuning is challenging. Most existing approaches rely on either manually annotated interaction trajectories (liu2024visualagentbench) that are high-quality but costly to obtain, or synthetic trajectories generated from large open-source or proprietary LLMs/VLMs, which may lack diversity and accuracy (qin2023toolllm; liu2024apigen; gao2024multi). Desirable GUI agents need to perform actions across a wide variety of applications, each with unique interaction sequences, while collecting comprehensive and diverse data across different applications and operating systems remains a major challenge. As a result, the lack of large-scale, diverse, and well-structured GUI trajectory data continues to be a key bottleneck in developing robust and generalized GUI agents.
@@
In this paper, we propose the TongUI framework that transforms millions of multimodal web tutorials into GUI trajectories for generalized GUI agents, where ‘Tong’ means generalization in Chinese. Our key observation is that there are readily available multimodal web tutorials on the Internet about how to control computers and smart mobile phones, offering a wealth of information on interacting with various applications and operating systems. These tutorials, in either videos or articles with screenshot formats, provide detailed step-by-step instructions on interacting with GUI. Compared to the aforementioned manually collected or synthetic data pipeline, online tutorials offer easy accessibility, rich information content, and good quality. Thus, it is a natural idea to convert the diverse multimodal web tutorials into task-solving trajectories for GUI agent tuning.
@@
In doing so, we design the TongUI framework composed of four steps: tutorial crawling, tutorial processing, trajectory generation, and data filtering. In the tutorial crawling step, we write some seed tasks and use LLMs to extend them into a wider collection of tasks. The generated tasks serve as keywords for retrieving content from hosts for online GUI tutorials (such as articles from WikiHow and videos from YouTube). The tutorial processing step aims to extract textual descriptions and screenshots of multimodal tutorials. We first obtain textual descriptions of multimodal tutorials via automatic speech recognition (ASR) or captioning, through which task queries and plans are produced using LLMs on the obtained textual descriptions. Then, we extract salient frames from videos as screenshots of each step, while the images in articles are directly regarded as screenshots. In the trajectory generation step, we leverage a zero-shot GUI agent to automatically recognize trajectories, including reasoning thoughts and actions between two steps. In the data filtering step, we apply a multi-stage pipeline to ensure data quality and GUI relevance, including duplicate tutorial removal, LLM-based content filtering, and trajectory-level filtering using GUI agents. Based on the TongUI framework, we construct a GUI-Net-1M dataset that contains 1M trajectories across five operating systems with more than 280 applications. As far as we know, GUI-Net-1M is the biggest open-source GUI trajectory dataset. Notably, our dataset contains tutorials from different time periods, where the same application may have evolving layouts and task-solving solutions, helping the model generalize to diverse GUI environments.
@@
In summary, our contributions are three-fold: (1) We construct GUI-Net-1M, having 1M trajectories across five operating systems and over 280 applications. As far as we know, it is the biggest open-source GUI trajectory dataset. (2) We propose the TongUI framework that enables GUI agents to automatically learn from rich web resources, leading to better generalization. (3) We develop the TongUI agent by fine-tuning Qwen2.5-VL models on GUI-Net-1M, achieving improvements on multiple popular benchmarks.
Related Work
GUI Agent
@@
The advancements of LLMs and VLMs accelerate the development of GUI agents. In the early state, only closed-source models are used (e.g., GPT-4, claude 3.5, and ChatGLM). Recently, more and more open-source models are released, such as Show-UI (lin2024showui), UI-TARS (qin2025ui), and CogVLM (wang2023cogvlm). Grounding and planning are two important capabilities in GUI agents. The grounding capability means whether the GUI could identify correct buttons or regions for operations. Some methods improve the grounding capability by adding more prompts (accessibility trees, HTML, and set-of-mask) (zhou2023webarena; deng2023mind2web; gur2023real; zhang2024ufo; lu2024omniparser) and collecting grounding data for fine-tuning (yang2024aria). As for the planning capability of GUI agents, existing efforts improve it by prompt engineering (jia2024agentstore; agashe2024agent), supervised fine-tuning (lin2024showui; xu2024aguvis; hong2024cogagent; qin2025ui), or reinforcement learning (lai2024autowebglm; putta2024agent). In addition, some efforts focus on evaluating the GUI agents from multiple aspects, including action sequence (rawles2024androidinthewild; he2024webvoyager), grounding precision (cheng2024seeclick; liu2024visualwebbench), trajectory effectiveness (lin2024videogui; li2025effects), and task completion (rawles2024androidworld; trivedi2024appworld; xie2025osworld). The evaluation performance in turn guides the research of intelligent agents.
Agent Data Collection
@@
Compared with data collection for LLMs or VLMs, collecting high-quality data for agents is more challenging, since agents usually require long trajectories that solve complex tasks in different domains (e.g., GUI, multimodal reasoning, embodied AI). Commonly used schemes include human annotation (liu2024visualagentbench; wang2025opencuaopenfoundationscomputeruse) and model synthesis (qin2023toolllm; liu2024apigen; gao2024multi). including the planning, action sequence, and action position of GUI agents (deng2023mind2web; rawles2024androidinthewild; chen2024guicourse; qin2025ui; cheng2024seeclick; lu2024gui). Note that the recent state-of-the-art method UI-TARS (qin2025ui) has not released its data. In contrast, we have fully released our data.
@@
In addition to the two schemes, considering that there are abundant resources for GUI operations on the Internet, some efforts focus on collecting trajectory data of GUI agents from the Web, to preserve the data quality and reduce costs simultaneously (ou2024synatra; xu2024agenttrek). They collect textual tutorials from the Internet, and convert them into trajectories by using LLMs or exploring in simulators. Unlike them, we directly collect multimodal tutorials from the Internet and propose a multimodal tutorial process pipeline to convert them into trajectories without any simulators. Meanwhile, compared with synthetic data in simulators, our data is closer to tasks in the real world with practical purposes, benefiting to the generalization capability of GUI agents. In table 1, we show the comparisons of GUI-Net-1M with existing GUI trajectory datasets, including GUI Odessey (lu2024gui), AgentTrek (xu2024agenttrek), AndroidControl (li2025effects), AGUVIS (xu2024aguvis), LBI (su2025learn), E-ANT (wang2024ant), ShowUI (lin2024showui), and AITW (rawles2023androidinthewild), where GUI-Net-1M has obvious advantages in data size, platform, and operating system numbers.
Datasets Size Platform OS
GUI Odyssey 7K W+M A
AgentTrek 10.4K W A+W+L+M+I
AndroidControl 15.3K W+M A
AGUVIS 35K W+M A+W+L
LBI 42.6K D+W L
E-ANT 49K M A
ShowUI 137K W+M W+L+I
AITW 715K W+M A
GUI-Net-1M 1M D+W+M A+W+L+M+I
Table 1: Comparison of GUI-Net-1M with other GUI trajectory datasets. For platform, ‘D’, ‘W’, and ‘M’ mean the desktop, web, and mobile, respectively. For operating systems (OS), ‘A, ‘W’, ‘L’, ‘M’, and ‘I’ denote Android, Windows, Linux, MacOS, and iOS, respectively.
@@
Method
Formulation
We formulate GUI tasks as a sequential decision process, and we adopt ReAct (yao2023react) as our agent framework. At one time step $i$, given the observation of previous $n$ steps $o_{i-n},\cdots,o_{i-1},o_{i}$ (e.g., a screenshot of GUI), the thoughts $r$ and actions $a$ of previous $n$ steps $(r_{i-n},a_{i-n},\cdots,r_{i-1},a_{i-1})$, and the query $q$, the agent generates a new thought $r_{i}$ and an executable action $a_{i}$ from the action space, such as clicking on a specific UI element, entering text, or scrolling through the interface. Then, executing the action $a_{i}$ leads to a new observation $o_{i+1}$ (e.g., an updated screenshot). This interaction loop continues, with the agent repeatedly observing the environment, selecting actions, and receiving updated observations until either a termination condition is met (e.g., task completes or fails) or the maximum number of steps is reached. We parameterize the GUI agent using a vision-language model $M_{\theta}$,
$r_{i}^{\star},a_{i}^{\star}$ $=\arg\max_{r_{i},a_{i}}M_{\theta}(r_{i},a_{i}\mid q,o_{i-n},r_{i-n},a_{i-n},$
$\quad\quad\quad\quad\quad\quad\ldots,o_{i-1},r_{i-1},a_{i-1},o_{i}).$ (1)
We propose the TongUI framework that aims to tune $M_{\theta}$ by learning from multimodal web tutorials, as shown in fig. 1. Such tutorials are usually formatted as $\{v,e\}$, where $v$ denotes the visual information (such as images or videos), and $e$ denotes the textual information (such as an introduction to a video). Our goal is to collect multimodal tutorials $\{v,e\}$, and convert them into training data $\{q,o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T}\}$ with the query $q$ and the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps.
Tutorial Source
To collect multimodal web tutorials, we carefully select multiple data sources via brainstorming, ensuring that we cover a wide range of applications, operating systems, and task types. Concretely, we choose YouTube, Bilibili, Wikihow, and Baidu Experience, which host much user-generated content on GUI tasks. YouTube and Bilibili have a variety of videos about GUI tasks. For video tutorials, the visual information $v$ is the video itself, and the textual information $e$ includes the title, the brief introduction to the video, and the caption or audio of the video. Wikihow and Baidu Experience are two widely used platforms that provide step-by-step articles with images in each step across applications and operating systems. Here, the visual information $v$ is the image sequence among all steps, and the textual information $e$ includes the title and textual content in the article. In this case, we could get diverse multimodal web tutorials.
Tutorial Crawling
We use a keyword-based search approach. The search keywords are constructed by combining the name of the applications or website with the task content, that is “app/web + task”. Here, “app/web” refers to the operating system or application being used (e.g., a mobile app or website), while “task” has a specific objective (e.g., changing font size in Word, or browsing in Chrome). We identify task seeds across multiple applications via brainstorming, and expanded them as search keywords using LLMs for more diverse and relevant tasks.
@@
Tutorial Process
Textual processing. Given a crawled multimodal tutorial $\{v,e\}$, this process extracts the task $q$ and rough descriptions $\{h_{1},\cdots,h_{T}\}$ in $T$ steps. For video tutorials from Youtube, Tiktok, and Bilibili, the task guidance is usually in the video’s audio information without any textual introductions. To solve this issue, we begin by applying the open-source speech recognition model, Whisper (radford2022robustsr), to transcribe the audio streams into the text information $e$. For image-text articles on Wikihow and Baidu Experience, we parse websites based on their article structures for the textual information $e$.
We use LLMs to identify key task-related verbs and nouns, such as “clicking a button”, “filling out a form”, or “selecting a menu”. Then, we leverage both the title and extracted text to classify tutorials into three categories: mobile, desktop, and others. The data with category “others” is discarded. Finally, we use LLMs to extract the task $q$ and summarize this process into rough descriptions $\{h_{1},\cdots,h_{T}\}$ of $T$ steps from the filtered textual information.
Visual processing. This process extracts the observation $\{o_{1},\cdots,o_{T}\}$ in the $T$ steps. For text-image articles, we could directly obtain the image sequences as the observation $\{o_{1},\cdots,o_{T}\}$ from the crawled data. In practice, there are some noisy data in Baidu Experience and Wikihow, which might be diagrams, comics, or natural images rather than GUI screenshots. To address this issue, we prompt GPT-4o-mini to classify each image as a screenshot or not.
For video tutorials, we need to extract key frames that correspond to critical actions, representing meaningful moments in the task-solving process. We observe that the audio transcript of the videos often contains valuable planning information, such as the description of task steps. Therefore, when audio is available, we first segment the video based on the audio transcript timestamps, treating each segment as a potential task step. We then extract key frames from each of these segments. In the absence of audio, we treat the entire video as a single segment and extract key frames throughout the video. In doing so, we adopt the MOG2 algorithm (zivkovic2004; zivkovic2004recursive; zivkovic2006) to detect significant changes.
@@
After textual and visual processing, tutorials from different sources are structured into a task $q$ with text and image pairs of $T$ steps, denoted as $\big(\{o_{1},h_{1}\},\cdots,\{o_{i},h_{i}\},\cdots,\{o_{T},h_{T}\}\big)$, where $o_{i}$ is the image and $h_{i}$ is the rough description in the $i$-th step.
Trajectory Generation
@@
This process aims to generate the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps using the task $q$, and image $o_{i}$ and rough description $h_{i}$ in $T$ steps. Concretely, we use a pretrained GUI agent (such as UI-TARS (qin2025ui)) to generate the thought $r_{i}$ and the action $a_{i}$ using the same action space defined by the GUI agent. For the thought $r_{i}$ and action $a_{i}$, we feed the observation $o_{i}$ with $h_{i}$ as the query to the zero-shot agent. Here, we use $h_{i}$ as the query for agents instead of using the task $q$, because $q$ usually contains an abstract goal instead of specific instructions about what to do on the observation. We empirically find that using $h_{i}$ leads to better performance. In this case, each step of text and image pairs can be written as $\{o_{i},r_{i},a_{i}\}$. In practice, the model might fail to generate well-formatted actions. We discard this step and make the next action the beginning of a new trajectory. For example, for a 4-step trajectory which has a failed generation in the second step, we split it into two training trajectories: $(\{o_{1},r_{1},a_{1}\})$ and $(\{o_{3},r_{3},a_{3}\},\{o_{4},r_{4},a_{4}\})$. Finally, we combine the trajectory with the task $q$ as $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$. We collect these trajectories into a GUI-Net-1M dataset that contains 1M trajectories. The data is on five operating systems with more than 280 applications.
Data Filtering
@@
We apply a multi-stage data filtering pipeline for data quality, as shown in fig. 1. (1) Duplicate Tutorial Removal. To eliminate redundancy of crawling, we directly remove exact duplicate tutorials based on their unique identifiers (e.g., video IDs and URLs). This step ensures that each retained tutorial is unique. (2) Content Filtering. After tutorial processing, we employ an LLM to semantically filter out tutorials that are irrelevant to the GUI tasks. Concretely, based on the article content, tutorial titles, and audio transcriptions, the LLM evaluates whether the tutorial is about GUI interactions. (3) Trajectory Filtering. In the trajectory generation stage, since we use UI-TARS (qin2025ui), if a screenshot corresponds to a step unrelated to GUI interaction, the agent typically predicts the action wait or call_user. We use this as a signal to discard such observations. Then, we feed the screenshot and trajectory of a task into a Qwen2.5-VL-7B model, and prompt it to filter out low-quality data.
Agent Tuning
@@
Given a data point $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$ of $T$-step, we train a VLM $M_{\theta}$ via supervised fine-tuning (SFT),
$\min\;\mathbb{E}_{(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})\sim\mathbb{D}}\Big[-\sum_{i=1}^{T}M_{\theta}(r_{i},a_{i}\mid q,o_{i-n},r_{i-n},$
$a_{i-n},\ldots,o_{i-1},r_{i-1},a_{i-1},o_{i})\Big]$ (2)